OpenStudy (anonymous):

Solve 5x + 14 - 2x = 9 - (4x + 2) x =3/7 x = - 1 x = 6 x = 8

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

3x+14=7-4x 7x=-7 x=-1